Literal Equations 1: Please answer showing your solutions.

{7} = 3/9 (x+5)

C=2/5 (x-3)
Answer #1

Soin 7-3(x+5) and c= C=3(2-3) 7= 9 63=3atis 48=32 .. n = 16 and, cs12-3) C = ? (16-3) 3 c - =16 and 2 x 13 S ...c = 26 S 26 =